OpenStudy (anonymous):

Square root of x+7 =5

OpenStudy (goten77):

\[\sqrt{x+7} = 5\]

OpenStudy (anonymous):

yes

OpenStudy (goten77):

\[x + 7 = 5^{2}\]

OpenStudy (goten77):

\[ x = 25 - 7\]

OpenStudy (goten77):

\[x = 18\]